### v2

# On quantum ergodicity for higher dimensional cat maps modulo prime powers

Abstract.
A discrete model of quantum ergodicity of linear maps generated by symplectic matrices $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$ modulo an integer $N\geqslant 1$, has been studied for $d=1$ and almost all $N$ by P. Kurlberg and Z. Rudnick (2001). Their result has been strengthened by J. Bourgain (2005) and then by A. Ostafe, I. E. Shparlinski and J. F. Voloch (2023). For arbitrary $d$ this has been studied by P. Kurlberg, A. Ostafe, Z. Rudnick and I. E. Shparlinski (2024). The corresponding equidistribution results, for certain eigenfunctions, share the same feature: they apply to almost all moduli $N$ and are unable to provide an explicit construction of such “good” values of $N$. Here, using a bound of I. E. Shparlinski (1978) on exponential sums with linear recurrence sequences modulo a power of a fixed prime, we construct such an explicit sequence of $N$, with a power saving on the discrepancy.
Key words and phrases:
Quantum unique ergodicity, linear map, exponential sums, prime powers
2020 Mathematics Subject Classification:
11L07, 81Q50
1. Introduction
1.1. Quantised linear maps and discrepancy of eigenfunctions
In what follows we freely borrow from the exposition in [13]. Namely, investigate equidistribution of eigenfunctions of the quantised cat map [6].
We need to introduce some notations.
For an integer $N\geq 1$ we denote by ${\mathbb{Z}}_{N}$ the residue ring modulo $N$ and consider the Hilbert space ${{\mathcal{H}}}_{N}=L^{2}\left(({\mathbb{Z}}_{N})^{d}\right)$ equipped with the scalar product
$\langle\varphi_{1},\varphi_{2}\rangle=\frac{1}{N^{d}}\sum_{\mathbf{u}\in{\mathbb{Z}}_{N}^{d}}\varphi_{1}(\mathbf{u})\overline{\varphi_{2}(\mathbf{u})},\qquad\varphi_{1},\varphi_{2}\in{{\mathcal{H}}}_{N}.$
In particular, the norm of $\varphi\in{{\mathcal{H}}}_{N}$ is given by
$\|\varphi\|=\langle\varphi,\varphi\rangle.$
We then consider the family of unitary operators
$\operatorname{T}_{N}(\mathbf{u}):{{\mathcal{H}}}_{N}\to{{\mathcal{H}}}_{N},\qquad\mathbf{u}=(\mathbf{u}_{1},\mathbf{u}_{2})\in{\mathbb{Z}}^{d}\times{\mathbb{Z}}^{d}={\mathbb{Z}}^{2d},$
which are defined by the following action on $\varphi\in{{\mathcal{H}}}_{N}$
(1.1) $\left(\operatorname{T}_{N}(\mathbf{u})\varphi\right)(\mathbf{w})={\mathbf{\,e}}_{2N}(\mathbf{u}_{1}\cdot\mathbf{u}_{2}){\mathbf{\,e}}_{N}(\mathbf{u}_{2}\cdot\mathbf{w})\varphi(\mathbf{w}+\mathbf{u}_{1}),$
[…]
$\operatorname{T}_{N}(\mathbf{u})\operatorname{T}_{N}(\mathbf{v})={\mathbf{\,e}}_{2N}\left(\omega\left(\mathbf{u},\mathbf{v}\right)\right)\operatorname{T}_{N}(\mathbf{u}+\mathbf{v}),$
where for $\mathbf{x}=(\mathbf{x}_{1},\mathbf{x}_{2})$, $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2})\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ we define
(1.2) $\omega(\mathbf{x},\mathbf{y})=\mathbf{x}_{1}\cdot\mathbf{y}_{2}-\mathbf{x}_{2}\cdot\mathbf{y}_{1},$
and
${\mathbf{\,e}}(z)=\exp\left(2\pi iz\right),\quad{\mathbf{\,e}}_{k}(z)={\mathbf{\,e}}(z/k),$
see also [16, Equation (2.6)].
For each real-valued function $f\in C^{\infty}(\mathbb{T}^{2d})$ (an “observable”), where $\mathbb{T}={\mathbb{R}}/Z$ is a unit torus, one associates a self-adjoint operator $\operatorname{Op}_{N}(f)$ on ${{\mathcal{H}}}_{N}$, analogous to a pseudo-differential operator with symbol $f$, defined by
(1.3) $\operatorname{Op}_{N}(f)=\sum_{\mathbf{u}\in{\mathbb{Z}}^{2d}}\widehat{f}(\mathbf{u})\operatorname{T}_{N}(\mathbf{u}),$
where
$f(\mathbf{x})=\sum_{\mathbf{u}\in{\mathbb{Z}}^{2d}}\widehat{f}(\mathbf{u}){\mathbf{\,e}}(\mathbf{u}\cdot\mathbf{x}).$
Denote by $\operatorname{Sp}(2d,{\mathbb{Z}})$ the group of all integer symplectic matrices $A$ which preserve the symplectic form (1.2), that is, $\omega(A\mathbf{x},A\mathbf{y})=\omega(\mathbf{x},\mathbf{y})$.
Associated to any $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$ is a quantum mechanical system. We briefly recall the key definitions:
Assuming $A\equiv I_{2d}\pmod{2}$, where $I_{2d}$ is the $2d$-dimensional identity matrix. For each $N\geq 1$, there is a unitary operator $U_{N}(A)$ on ${{\mathcal{H}}}_{N}$ such that that for every $f\in C^{\infty}(\mathbb{T}^{2d})$, we have the exact Egorov property
$U_{N}(A)^{*}\operatorname{Op}_{N}(f)U_{N}(A)=\operatorname{Op}_{N}(f\circ A),$
where $U_{N}(A)^{*}=\overline{U_{N}(A)}^{\,t}$, we refer to [5, 14, 15, 16, 17, 22] for a detailed exposition in the case $d=1$ and [9] for higher dimensions.
We further assume that $A$ has an irreducible characteristic polynomial (and thus is diagonalisable over $\mathbb{C}$}) and there are no roots of unity amongst the eigenvalues of $A$ and their nontrivial ratios.
General results of the Quantum Ergodicity Theorem [24, 26, 3], make it natural to expect that any normalised sequence of eigenfunctions $\psi_{N}\in{{\mathcal{H}}}_{N}$ of the operator $U_{N}(A)$ satisfy
$\lim_{N\to\infty}\left\langle\operatorname{Op}_{N}(f)\psi_{N},\psi_{N}\right\rangle=\int_{\mathbb{T}^{2d}}f(\mathbf{x})d\mathbf{x}$
for all $f\in C^{\infty}(\mathbb{T}^{2d})$, in which case we say that the sequence of eigenfunctions $\{\psi_{N}\}$ is uniformly distributed, see also [23].
To make this more quantitative, we introduce the following definition of the discrepancy
$\Delta_{A}(f,N)=\max_{\psi\in\varPsi_{N}(A)}\left|\left\langle\operatorname{Op}_{N}(f)\psi,\psi\right\rangle-\int_{\mathbb{T}^{2d}}f(\mathbf{x})d\mathbf{x}\right|,$
where $\varPsi_{N}(A)$ is the set of all normalised (that is, with $\|\psi\|=1$) eigenfunctions $\psi$ of $U_{N}(A)$ in ${{\mathcal{H}}}_{N}$.
Remark 1.1.
We note that the notion of discrepancy is sometimes called, especially in mathematical physics literature, the rate of decay of matrix coefficients.
Then the uniformity of distribution property for $N$ running through a certain infinite sequence $\mathcal{N}$$\subseteq{\mathbb{N}}$ means that we ask if for all $f\in C^{\infty}(\mathbb{T}^{2d})$,
(1.4) $\lim_{\begin{subarray}{c}N\to\infty\\ N\in{\mathcal{N}}\end{subarray}}\Delta_{A}(f,N)=0.$
In turn this leads to the following:
Problem 1.2.
Make the class of sequences $\mathcal{N}$} for which (1.4) holds as broad as possible.
Problem 1.2 has first been addressed in the work of Kurlberg and Rudnick [16] where (1.4), for $d=1$, has been established for almost all $N$, that is. when $\mathcal{N}$} is a set of asymptotic density $1$. Bourgain [1] has used methods of additive combinatorics to give a bound with a power saving $\Delta_{A}(f,N)\leqslant N^{-\delta}$, for some unspecified $\delta>0$ and also for almost all $N$. Finally, using a different approach via methods and results of algebraic geometry, Ostafe, Shparlinski and Voloch [20] have shown that one can take any $\delta<1/60$ in the above bound. For $d\geqslant 2$, the only known result is due to Kurlberg, Ostafe, Rudnick and Shparlinski [13], which gives (1.4) in any dimension. We remark that although the approaches in [16] and [13] are able to produce and explicit bound on the rate of convergence in (1.4), they are incapable of giving a power saving. We also note recent works [4, 10, 21] of somewhat different flavour.
Here concentrate on a different aspect of this question and address the following:
Problem 1.3.
Construct an explicit sequence $\mathcal{N}$}, which admit strong bounds, preferably with a power saving, on the rate of convergence in (1.4).
1.2. Construction and the discrepancy bound
Below, we always assume that $A\equiv I_{2d}\pmod{2}$ and that the characteristic polynomial $f_{A}$ of the matrix $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$ is irreducible over $\mathbb{Z}$}. In particular, $A$ is diagonalisable over $\mathbb{C}$}. We also assume that there are no roots of unity amongst the eigenvalues of $A$ and their nontrivial ratios.
Let $p>2d$ be a fixed prime such that $f_{A}$ splits completely modulo $p$ and has $2d$ distinct roots (that is $p$ does not divide the discriminant of $f_{A}$). We additionally assume that $p\nmid\det A$. By the Chebotarev Density Theorem, the set of such primes $p$ is of positive relative density in the set of all primes.
Our sequence of “good” moduli, required for Problem 1.3 is simply the sequence of powers
(1.5) ${\mathcal{N}}=\{p^{k}:\leavevmode\nobreak\ k=0,1,\ldots\}.$
Our main result establishes (1.4) for the above sequence $\mathcal{N}$} with a reasonably strong bound on rate of convergence.
Let
[…]
We note that for $d\geqslant 2$ we have $\kappa_{d}\geqslant 1/(4d-1)$.
Theorem 1.4.
For the sequence $\mathcal{N}$} given by (1.5) and $N\in{\mathcal{N}}$, we have
$\Delta_{A}(f,N)\leqslant N^{-\kappa_{d}+o(1)}$
as $N\to\infty$.
The proof is based on a link between between $\Delta_{A}(f,N)$ and bounds on the number of solutions on certain systems of congruences, first established in [16] and then generalised and used in all other papers on this subject [1, 13, 20]. In turn, we estimate the aforementioned number of solutions, using bounds of exponential sums with linear recurrence sequences from [25].
We note that the cat map modulo prime powers has also been studied by Kelmer [8] and Olofsson [19], but their results are of different flavour.
1.3. Notation
Throughout the paper, the notations
$X=O(Y),\qquad X\ll Y,\qquad Y\gg X$
are all equivalent to the statement that the inequality $|X|\leqslant cY$ holds with some constant $c>0$, which may depend on the matrix $A$.
We recall that the additive character with period $1$ is denoted by
$z\in\mathbb{R}\ \mapsto\ {\mathbf{\,e}}(z)=\exp\left(2\pi iz\right).$
For an integer $q\geqslant 1$ it is also convenient to define
${\mathbf{\,e}}_{q}(z)={\mathbf{\,e}}(z/q).$
The letter $p$, with or without indices, always denotes prime numbers.
Given an algebraic number $\gamma$ we denote by $\operatorname{ord}(\gamma,N)$ its order modulo $N$ (assuming that the ideals generated by $\gamma$ and $N$ are relatively prime in an appropriate number field). In particular, for an element $\lambda\in{\mathbb{F}}_{p^{s}}$, $\operatorname{ord}(\lambda,p)$ represents the order of $\lambda$ in ${\mathbb{F}}_{p^{s}}$.
Similarly, we use $\operatorname{ord}(A,N)$ to denote the order of $A$ modulo $N$ (which always exists if $\gcd(\det A,N)=1$ and in particular for $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$).
Finally, we use $\nu_{p}(z)$ to denote the $p$-adic order of $z\in{\mathbb{Q}}_{p}$, where ${\mathbb{Q}}_{p}$ is the field of $p$-adic numbers.
2. Operators $\operatorname{T}_{N}$ and congruences
2.1. Preliminaries
As in [13], and then also in [1, 13, 20], we observe that it is enough bound the quantity $\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle$, where
$\operatorname{T}_{N}(\mathbf{u})=\operatorname{Op}_{N}({\mathbf{\,e}}\left(\mathbf{x}\cdot\mathbf{u}\right)),$
see also (1.3)), and $\psi\in\Psi_{N}(A)$ runs through eigenfunctions of $U_{N}(A)$, with frequency $\mathbf{u}$ growing slowly with $N$ (for example, as any power $N^{\eta}$ for any fixed $\eta>0$).
We also use $\operatorname{ord}(A,N)$ to denote the order of $A$ modulo $N$, which is always correctly defined if $\gcd(\det A,N)=1$, which we always assume.
For a row vector $\mathbf{u}\in{\mathbb{Z}}^{2d}$, $\mathbf{u}\not\equiv{\mathbf{0}}_{2d}\pmod{N}$, where ${\mathbf{0}}_{2d}$ is the $2d$-dimensional zero-vector, we denote by $Q_{s}(N;\mathbf{u})$ the number of solutions of the congruence
(2.1) $\mathbf{u}\left(A^{x_{1}}+\ldots+A^{x_{s}}-A^{y_{1}}-\ldots-A^{y_{s}}\right)\equiv{\mathbf{0}}_{2d}\pmod{N}$,
with integers $1\leqslant x_{i},y_{i}\leqslant\operatorname{ord}(A,N)$, $i=1,\ldots,s$.
The key inequality below connects the $2s$-th moment associated to the basic observables $\operatorname{T}_{N}(\mathbf{u})$ with the number of solutions $Q_{s}(N;\mathbf{u})$ to the system (2.1). For even $s$, this is given (in broader generality) by [13, Lemma 4.1]. However this parity condition is too restrictive for us, hence we show how to prove a result for any $s$.
Lemma 2.1.
Let $\mathbf{u}\in{\mathbb{Z}}^{2d}\setminus\{{\mathbf{0}}_{2d}\}$. Then
$\max_{\psi\in\Psi_{N}(A)}\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|^{2s}\leqslant N^{d}\frac{Q_{s}(N;\mathbf{u})}{\operatorname{ord}(A,N)^{2s}},$
where the maximum is taken over all normalised eigenfunctions of $U_{N}(A)$.
Proof.
We argue exactly as in the proof of [13, Lemma 4.1]. Denote $\tau=\operatorname{ord}(A,N)$, and consider
$D(\mathbf{u})=\frac{1}{\tau}\sum_{i=1}^{\tau}\operatorname{T}_{N}(\mathbf{u}A^{i}),\quad\mathrm{and}\quad H(\mathbf{u})=D(\mathbf{u})^{*}D(\mathbf{u}).$
We have
$|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle|^{2s}\leqslant\|D(\mathbf{u})\|^{2s}=\|H(\mathbf{u})\|^{s}=\|H(\mathbf{u})^{s}\|,$
where $\|\cdot\|$ denotes the operator norm.
At this point, our argument differs from the proof of [13, Lemma 4.1]. We note that $H(\mathbf{u})$ is not only Hermitian, but also a positive semidefinite matrix; this is because
$\mathbf{z}^{*}H(\mathbf{u})\mathbf{z}=\mathbf{z}^{*}D(\mathbf{u})^{*}D(\mathbf{u})\mathbf{z}=\|D(\mathbf{u})\mathbf{z}\|^{2}.$
Moreover, the operator $H(\mathbf{u})$ is also unitarily diagonalisable (see [7, Theorem 2.5.6]), with non-negative eigenvalues. This shows that $H(\mathbf{u})^{s}$ is also a positive semidefinite matrix. Now, it is not hard to see that $\|H(\mathbf{u})^{s}\|=\rho(H(\mathbf{u})^{s})$, where $\rho(H(\mathbf{u})^{s})$ is the spectral radius of $H(\mathbf{u})^{s}$, that is, the maximum of all the eigenvalues of $H(\mathbf{u})^{s}$. Clearly, we then have
$\|H(\mathbf{u})^{s}\|\leqslant\operatorname{tr}(H(\mathbf{u})^{s}).$
Now the proof concludes, by the same treatment as in the proof of [13, Lemma 4.1]. ∎
Next we reduce $Q_{s}(N;\mathbf{u})$ to the number of solutions to a similar system of equations but without the vector $\mathbf{u}$.
2.2. Linear independence of matrix powers
For a vector $\mathbf{z}=(z_{1},\ldots,z_{n})\in{\mathbb{R}}^{n}$, as usual, we denote
$\|\mathbf{z}\|_{2}=(z_{1}^{2}+\ldots+z_{n}^{2})^{1/2}.$
We need the following result which is given by  [13, Lemma 4.3, (i)] (in broader generality).
Throughout this section we always assume that $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$ has an irreducible characteristic polynomial.
Lemma 2.2.
For any non-zero row vector $\mathbf{u}\in{\mathbb{Z}}^{2d}$, the vectors
$\mathbf{u},\mathbf{u}A,\ldots,\mathbf{u}A^{2d-1}$
are linearly independent.
We are now ready to establish the desired result which allows to remove $\mathbf{u}$ in our considerations of $Q_{s}(N;\mathbf{u})$.
Lemma 2.3.
There is a constant $C(A)$ depending only on $A$, such that if we have $p^{m+1}>C(A)\|\mathbf{u}\|_{2}^{2d}$ for some integer $1\leqslant m<k$, then for any solution $\left(x_{1},\ldots,x_{s},y_{1},\ldots,y_{s}\right)\in{\mathbb{Z}}^{2s}$ to (2.1) with $N=p^{k}$, we have
$A^{x_{1}}+\ldots A^{x_{s}}\equiv A^{y_{1}}+\ldots+A^{y_{s}}\pmod{p^{k-m}}$.
Proof.
Let us set
$B=A^{x_{1}}+\ldots A^{x_{s}}-A^{y_{1}}-\ldots-A^{y_{s}}.$
Since $\mathbf{u}B\equiv 0\pmod{p^{k}}$, considering the matrix $X$ whose rows are $\mathbf{u},\mathbf{u}A,\ldots,\mathbf{u}A^{2d-1}$ and observing that $A$ and $B$ commute, we have $XB\equiv 0\pmod{p^{k}}$. In particular, multiplying both sides by the adjoint of $X$, we get
(2.2) $\det X\cdot B\equiv 0\pmod{p^{k}}$.
On the other hand, Lemma 2.2 shows that $\det X$ is a non-zero integer. In particular, if $p^{m+1}\nmid\det X$, then the congruence (2.2) implies that $B\equiv 0\pmod{p^{k-m}}$. The proof now follows, as we obviously have $\det X\ll\|\mathbf{u}\|_{2}^{2d}$. ∎
Let $p$ be a split prime which does not divide the discriminant and the constant coefficient of the characteristic polynomial of $A$ (that is, exactly as we assume in Section 1.2).
We see that we have $2d$ distinct the eigenvalues of $A$ modulo $p$, that is, in the finite field ${\mathbb{F}}_{p}$ of $p$ elements, which using Hensel lifting give us the roots
$\lambda_{1},\ldots,\lambda_{2d}\in{\mathbb{Z}}/p^{k}{\mathbb{Z}},$
of the characteristic polynomial of $A$ modulo $p^{k}$.
We have the following variant of [13, Lemma 4.4].
Lemma 2.4.
Let $p$ be any prime as in Section 1.2, and let $m$ be the smallest integer with $p^{m+1}>C(A)\|\mathbf{u}\|_{2}^{2d}$ where $C(A)$ is as in Lemma 2.3. For any solution $(x_{1},\ldots x_{s},y_{1},\ldots,y_{s})$ to (2.1) with $N=p^{k}$ and $k>m$, we have
$\lambda_{i}^{x_{1}}+\ldots+\lambda_{i}^{x_{s}}\equiv\lambda_{i}^{y_{1}}+\ldots+\lambda_{i}^{y_{s}}\pmod{p^{k-m}},\quad i=1,\ldots,2d.$
Proof.
By the assumption on $p$, clearly the characteristic polynomial of $A$ has $2d$ distinct roots in ${\mathbb{Q}}_{p}$. In particular, $A$ is diagonalisable over ${\mathbb{Q}}_{p}$. Denote $\bm{\lambda}_{1},\ldots,\bm{\lambda}_{2d}\in{\mathbb{Z}}_{p}$ be its eigenvalues, where ${\mathbb{Z}}_{p}$ is the ring of $p$-adic integers in ${\mathbb{Q}}_{p}$. We have
(2.3) $\nu_{p}\left(\lambda_{i}-\bm{\lambda}_{i}\right)\geqslant k,\qquad i=1,\ldots,2d.$
For each $1\leqslant i\leqslant 2d$, there exists a non-zero vector $\mathbf{v}_{i}\in({\mathbb{Q}}_{p})^{2d}$, for which $\mathbf{v}_{i}A=\bm{\lambda}_{i}\mathbf{v}_{i}$. We scale $\mathbf{v}_{i}=\left(v_{i,1},\ldots,v_{i,2d}\right)$ so that all its coordinates lie in ${\mathbb{Z}}_{p}$, with some coordinate $v_{i,j_{i}}$ satisfying
(2.4) $\nu_{p}\left(v_{i,j_{i}}\right)=0.$
Lemma 2.3 then implies that
$\mathbf{v}_{i}(\bm{\lambda}_{i}^{x_{1}}+\ldots\bm{\lambda}_{i}^{x_{s}}-\bm{\lambda}_{i}^{y_{1}}-\ldots-\bm{\lambda}_{i}^{y_{s}})\in\left(p^{k-m}{\mathbb{Z}}_{p}\right)^{2d},\qquad i=1,\ldots,2d,$
and thus
$\nu_{p}\left(v_{i,j_{i}}(\bm{\lambda}_{i}^{x_{1}}+\ldots\bm{\lambda}_{i}^{x_{s}}-\bm{\lambda}_{i}^{y_{1}}-\ldots-\bm{\lambda}_{i}^{y_{s}}\right)\geqslant k-m,\qquad i=1,\ldots,2d.$
The result now follows from (2.3) and (2.4). ∎
Hence, we see from Lemmas 2.3 and 2.4, that
(2.5) $Q_{s}(p^{k};\mathbf{u})\ll p^{2sm}R_{s}(p^{k-m}),$
where $m$ is as in Lemma 2.3 and $R_{s}(p^{r};\mathbf{u})$ is the number of solutions to the system of equations
(2.6) $\lambda_{i}^{x_{1}}+\ldots+\lambda_{i}^{x_{s}}\equiv\lambda_{i}^{y_{1}}+\ldots+\lambda_{i}^{y_{s}}\pmod{p^{r}},\quad i=1,\ldots,2d,$
in variables $x_{1},y_{1},\ldots,x_{s},y_{s}=1,\ldots,\operatorname{ord}(A,p^{k})$.
3. Multiplicative orders and exponential sums
3.1. Multiplicative orders
We need to collect the simple and well-known properties of multiplicative orders modulo prime powers. More general results have been given by Korobov [11, 12], we need the following direct consequence of [11, Lemma 1].
Lemma 3.1.
Assume that a prime $p\geqslant 3$ and an integer $\lambda\neq\pm 1$ are relatively prime. Let
$\gamma=\nu_{p}\left(\lambda^{\operatorname{ord}(\lambda,p)}-1\right).$
Then for $k\geqslant\gamma$ we have
$\operatorname{ord}(\lambda,p^{k})=\operatorname{ord}(\lambda,p)p^{k-\gamma}.$
Define integers $\rho_{i,j}$, $1\leqslant i,j\leqslant 2d$, by $\rho_{i,j}\equiv\lambda_{i}/\lambda_{j}\pmod{p^{k}}$.
We also define $\gamma_{i}$ and $\gamma_{i,j}$ as in Lemma 3.1 for $\lambda_{i}$ and $\rho_{i,j}$, respectively, $1\leqslant i,j\leqslant 2d$
Lemma 3.2.
There is a constant $c(A,p)$ depending only on $A$ and $p$, such that
$\max_{\begin{subarray}{c}1\leqslant i,j\leqslant 2d\\ i\neq j\end{subarray}}\{\gamma_{i},\gamma_{i,j}\}\leqslant c(A,p).$
Proof.
Let $\mu_{1},\ldots,\mu_{2d}$ be the eigenvalues $A$ and $\mathfrak{p}$ be a prime ideal of ${\mathbb{Q}}\left(\mu_{1},\ldots,\mu_{2d}\right)$. Note that $p^{\gamma_{i}}\mid\lambda_{i}^{t}-1$, with integer $t\geqslant 1$ implies $\mathfrak{p}^{\gamma_{i}}\mid\mu_{i}^{t}-1$. By our assumption on $A$ we have $\mu_{i}^{t}-1\neq 0$ and since $\operatorname{ord}(\lambda,p)\leqslant p-1$ for any integer $\lambda\not\equiv 0\pmod{p}$, we see that $\gamma_{i}$ is bounded only in terms of $A$ and $p$.
Similarly, $p^{\gamma_{i,j}}\mid\rho_{i,j}^{t}-1$ implies $\mathfrak{p}^{\gamma_{i,j}}\mid\mu_{i}^{t}-\mu_{j}^{t}$ and for $i\neq j$ the same argument applies. ∎
3.2. Exponential sums
Let $p$ be any prime as in Section 1.2 and let $\lambda_{1},\ldots\lambda_{2d}$ be as in Section $2.$
For a vector of integers $\mathbf{a}=\left(a_{1},\ldots,a_{2d}\right)$ and a positive integer $r$, we define the exponential sums
$S_{r}(\mathbf{a})=\sum_{x=1}^{t_{r}}{\mathbf{\,e}}_{p^{r}}\left(a_{1}\lambda_{1}^{x}+\ldots+a_{2d}\lambda_{2d}^{x}\right),$
where $t_{r}$ is the period of the sequence $a_{1}\lambda_{1}^{x}+\ldots+a_{2d}\lambda_{2d}^{x}$, $x=0,1,\ldots$, modulo $p^{r}$.
We note that the following bound on these exponential sums is essentially established in [25] and in fact for essentially arbitrary linear recurrence sequences. Note that a similar argument has also been used in [18].
Lemma 3.3.
Let $p$ be any prime as in Section 1.2 and let $\lambda_{1},\ldots\lambda_{2d}$ be as in Section $2.$ Then for any integer $r\geqslant 1$, uniformly over integers $a_{1},\ldots,a_{2d}$ with
$\gcd(a_{1},\ldots,a_{2d},p)=1,$
we have
$\left|S_{r}(\mathbf{a})\right|\leqslant t_{r}^{1-1/(2d)+o(1)},\qquad\text{as}\ r\to\infty.$
Proof.
The result in [25, Theorem 2] is formulated for fixed integers $\lambda_{1},\ldots\lambda_{2d}$ (and fixed $d$ and $p$). Since the work [25] is difficult to access, we now summarise some ideas used in the proof, which references to much easier accessible work [18]. In full generality, [25, Theorem 2] gives the following bound
$\left|\sum_{x=1}^{\tau_{r}}{\mathbf{\,e}}_{p^{r}}\left(u(x)\right)\right|\leqslant\tau_{r}^{1-1/e+o(1)}$
on exponential sums over the full period $\tau_{r}$ modulo $p^{r}$ of an integer linear recurrence sequence $u(x)$ of order $e$, with a square-fee characteristic polynomial $f(X)\in{\mathbb{Z}}[X]$, such that there are no roots of unity amongst the roots of $f$ and their nontrivial ratios. This bound is based on:
  • a polynomial representation (as polynomials in $y$) of the sequences $u(a+\tau_{s}y)$ for each $a=0,\ldots,\tau_{s}-1$, with $s$ slowly growing with $r$, and an upper bound on $p$-adic order of at least one coefficient among every $e$ consecutive coefficients of this polynomial, see [18, Lemma 2.5];
  • a bound on exponential sums
    (3.1) $\sum_{y=1}^{p^{r}}{\mathbf{\,e}}_{p^{r}}\left(F(y)\right)\ll p^{r(1-1/e)}$
    provided $p>e$, with any polynomial $F(Y)\in{\mathbb{Z}}[Y]$ of the form $F(Y)=pG(Y)+A_{e}Y^{e}+\ldots+A_{1}Y$ for an arbitrary $G(Y)\in{\mathbb{Z}}[Y]$, and $\gcd(A_{1},\ldots,A_{e},p)=1$, which follows, for example, from a much more general result of Cochrane and Zheng [2, Theorem 3.1] (with the implied constant depending only on $e$ and $p$).
It is important to recall that the implied constant in (3.1) depends only on $e$ and $p$, in particular, it does not depend on $\deg F$.
Note that we have $e=2d$ in our setting.
As we have mentioned a similar strategy has also been used in [18], where instead of the above complete sums, very short exponential sums are used.
Examining the dependencies in implied constants throughout the argument of the proof of [25, Theorem 2] one can easily verify that in fact all constants depend only on $d$, $p$ and parameters $\gamma_{i}$ and $\gamma_{i,j}$ from Lemma 3.2, which depend only on the matrix AA and the prime pp. ∎
Remark 3.4.
Certainly the parity of the number of terms in the sums $S_{r}(\mathbf{a})$, plays no role in argument and the similar statement holds for any number $e$ of terms instead of $2d$ (with the saving $1/e$).
3.3. Bounding $R_{2}(p^{r})$
Here we use the idea of Kurlberg and Rudnick [16] to estimate $R_{2}(p^{r})$. While it is not necessary for getting a power saving on $\Delta_{A}(f,p^{k})$ in Theorem 1.4, it leads to a larger value of $\kappa_{d}$.
Lemma 3.5.
Let $p$ be any prime as in Section 1.2 and let $m$ and $\lambda_{1},\ldots\lambda_{2d}$ be as in Section $2.$ Then
$R_{2}(p^{r})\ll r^{2}p^{7r/3}.$
Proof.
Since $A\in\operatorname{Sp}(2d,{\mathbb{Z}})$, we can choose an arbitrary pair of eigenvalues of the form $(\lambda,\lambda^{-1})$ and use only two corresponding equations from the system (2.6). Hence, we consider the system
$\lambda^{x_{1}}+\lambda^{x_{2}}\equiv\lambda^{y_{1}}+\lambda^{y_{2}}\pmod{p^{r}}$,
$\lambda^{-x_{1}}+\lambda^{-x_{2}}\equiv\lambda^{-y_{1}}+\lambda^{-y_{2}}\pmod{p^{r}}$,
in variables $x_{1},x_{2},y_{1},y_{2}=1,\ldots,\tau$, where $\tau=\operatorname{ord}(\lambda,p^{r})$.
Denoting $u=x_{1}-y_{1}$ and $v=x_{2}-y_{1}$ and repeating the same argument as in the proof of [16, Lemma 5] we derive
$\left(1-\lambda^{u}\right)\left(1-\lambda^{v}\right)\left(\lambda^{u-v}+1\right)\equiv 0\pmod{p^{r}}$,
see [16, Equation (4.7)].
We now fix (in $\tau$ possible ways) the value of $y_{1}$.
Next we fix integers $\omega_{1},\omega_{2},\omega_{3}\geqslant 0$ with
$\omega_{1}+\omega_{2}+\omega_{3}=r$
and count pairs $(u,v)$ for which the corresponding $p$-adic orders satisfy
$\nu_{p}\left(1-\lambda^{u}\right)\geqslant\omega_{1},\quad\nu_{p}\left(1-\lambda^{v}\right)\geqslant\omega_{2},\quad\nu_{p}\left(\lambda^{u-v}+1\right)\geqslant\omega_{3}.$
We choose tWo largest values, say $\omega_{a}$ and $\omega_{b}$, $1\leqslant a<b\leqslant 3$ and note that we clearly have $\omega_{a}+\omega_{b}\geqslant 2r/3$.
Thus, using (3.1), we see that for a fixed (in $\tau$ possible ways) value of $y_{1}$, the pairs $(u,v)$ take at most $O\left(p^{4r/3}\right)$ values.
Indeed, without loss of generality we can assume, that $a=1$. Hence, for each fixed $y_{1}$ by Lemma 3.1, there are $O\left(p^{r-\omega_{a}}\right)$ values for $u$ and hence to $x_{1}$. We now see that whether $b=2$ or $b=3$, there are $O\left(p^{r-\omega_{b}}\right)$ values for $v$ and hence to $x_{2}$.
Hence, for each choice of $\omega_{1},\omega_{2},\omega_{3}$ we have
$O\left(\tau p^{4r/3}\right)=O\left(p^{7r/3}\right)$
choices for the triple $(x_{1},x_{2},y_{1})$, after which $y_{2}$ is uniquely defined.
Since there are at most $r^{2}$ possible choices for $\omega_{1},\omega_{2},\omega_{3}$, the desired bound follows. ∎
4. Proof of Theorem 1.4
4.1. Bounding $Q_{s}(N;\mathbf{u})$ via the fourth moment
Let $p$ be any prime as in Section 1.2 and let $N=p^{k}$. Assume that $k>m$ where $m$ is the smallest integer with $p^{m+1}>C(A)\|\mathbf{u}\|_{2}^{2d}$ where $C(A)$ is as in Lemma 2.3. Denote $T=\operatorname{ord}(A,N)=\operatorname{ord}(A,p^{k})$.
Using the orthogonality of exponential functions, it follows from (2.5) and (2.6) that
(4.1) $\begin{split}Q_{s}&(N;\mathbf{u})\\ &\leqslant\frac{1}{p^{2d(k-m)}}\sum_{\mathbf{a}\in({\mathbb{Z}}/p^{k-m}{\mathbb{Z}})^{2d}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{k-m}}\left(a_{1}\lambda_{1}^{x}+\ldots+a_{2d}\lambda_{2d}^{x}\right)\right|^{2s}.\end{split}$
For each $r=0,\ldots,k-m$ we separate the contribution
$W_{r}$ $=\sum_{\begin{subarray}{c}\mathbf{a}\in({\mathbb{Z}}/p^{k-m}{\mathbb{Z}})^{2d}\\ \gcd\left(a_{1},\ldots,a_{2d},p^{k-m}\right)=p^{k-m-r}\end{subarray}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{k-m}}\left(a_{1}\lambda_{1}^{x}+\ldots+a_{2d}\lambda_{2d}^{x}\right)\right|^{2s}$
$=\sum_{\begin{subarray}{c}\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}\\ \gcd\left(b_{1},\ldots,b_{2d},p\right)=1\end{subarray}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|^{2s}\$
to the sum on the right hand side of (4.1) from vectors $\mathbf{a}$ for which $\gcd\left(a_{1},\ldots,a_{2d},p^{k-m}\right)=p^{k-m-r}$.
For each $\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}$ with $\gcd\left(b_{1},\ldots,b_{2d},p\right)=1$ we see that the period $t_{r}(\mathbf{b})$ of the sequence $b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}$ modulo $p^{r}$ satisfies
$t_{r}(\mathbf{b})\gg p^{r}\qquad\mbox{and}\qquad t_{r}(\mathbf{b})\mid t_{k-m}(\mathbf{b})\mid T,$
and hence by Lemma 3.3 we have
(4.2) $\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|\ll T^{1+o(1)}p^{-r/(2d)}.$
Therefore, assuming that
(4.3) $s\geqslant 2$
and applying (4.2) 2s42s-4 times, we derive
$W_{r}$ $\leqslant\left(T^{1+o(1)}p^{-r/(2d)}\right)^{2s-4}$
$\qquad\qquad\times\sum_{\begin{subarray}{c}\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}\\ \gcd\left(b_{1},\ldots,b_{2d},p\right)=1\end{subarray}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|^{4}$
$\leqslant T^{2s-2+o(1)}p^{-r(s-2)/d}\sum_{\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|^{4}.$
It is easy to see that first by the orthogonality of exponential functions and then by Lemmas 3.1 and 3.5 we have
$\sum_{\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}}\left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|^{4}$ $\leqslant(T/t_{r})^{4}p^{2dr}R_{2}(p^{r})$
$\leqslant k^{2}T^{4}p^{2dr-5r/3+o(1)}.$
Hence,
$W_{r}\leqslant k^{2}T^{2s+o(1)}p^{r\left(2d-5/3-(s-2)/d\right)}.$
First we assume that
(4.4) $s-2\leqslant d(2d-5/3),$
and noting that $p^{m}\ll\|\mathbf{u}\|_{2}^{2d}$ and $k\ll\log N\ll\log T$, we derive from (4.1) that
$Q_{s}(N;\mathbf{u})$ $\leqslant\frac{1}{p^{2d(k-m)}}\sum_{r=0}^{k-m}W_{r}$
$\leqslant\frac{k^{3}}{p^{2d(k-m)}}T^{2s}p^{(k-m)\left(2d-5/3-(s-2)/d\right)}$
$\leqslant\|\mathbf{u}\|_{2}^{10d/3+2s-4}T^{2s+o(1)}p^{-k\left(5/3+(s-2)/d\right)}$
It now follows from Lemma 2.1 that,
(4.5) $\begin{split}\max_{\psi\in\Psi_{N}(A)}&\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|\\ &\leqslant\|\mathbf{u}\|_{2}^{(10d/3+2s-4)/(2s)}p^{k\left(5/3+(s-2)/d\right)/(4s)}T^{o(1)}\\ &\leqslant\|\mathbf{u}\|_{2}^{(s-2+2d)/s}N^{-\left(5/3+(s-2)/d-d\right)/(2s)+o(1)}.\end{split}$
Now we consider that case
(4.6) s2>d(2d5/3).s-2>d(2d-5/3).
In this case we obtain
$Q_{s}(N;\mathbf{u})\leqslant\frac{k^{2}}{p^{2kd}}\|\mathbf{u}\|_{2}^{4d^{2}}T^{2s+o(1)},$
which by Lemma 2.1 implies that
(4.7) $\max_{\psi\in\Psi_{N}(A)}\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|\leqslant\|\mathbf{u}\|_{2}^{2d^{2}/s}N^{-d/(2s)+o(1)}.$
4.2. Bounding $Q_{s}(N;\mathbf{u})$ via the second moment
We now establish yet another bound on $Q_{s}(N;\mathbf{u})$, and thus on $\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle$, which is better than (4.5) for $d=1,2$.
We proceed as before, but now we use (4.2) 2s22s-2 times and also use the orthogonality relation
$\sum_{\mathbf{b}\in({\mathbb{Z}}/p^{r}{\mathbb{Z}})^{2d}}\ \left|\sum_{x=1}^{T}{\mathbf{\,e}}_{p^{r}}\left(b_{1}\lambda_{1}^{x}+\ldots+b_{2d}\lambda_{2d}^{x}\right)\right|^{2}\leqslant p^{2dr}T(T/t_{r})\ll T^{2}p^{r(2d-1)},$
instead of Lemma 3.5. This time, assuming that
(4.8) $s-1\leqslant d(2d-1),$
we derive from (4.1) that
$Q_{s}(N;\mathbf{u})$ $\leqslant\frac{1}{p^{2d(k-m)}}\sum_{r=0}^{k-m}W_{r}$
$\leqslant\frac{k}{p^{2d(k-m)}}T^{2s+o(1)}p^{(k-m)\left(2d-1-(s-1)/d\right)}$
$\leqslant\|\mathbf{u}\|_{2}^{2s+2d-2}T^{2s+o(1)}p^{-k\left(1+(s-1)/d\right)}.$
It now follows from Lemma 2.1 that,
(4.9) $\begin{split}\max_{\psi\in\Psi_{N}(A)}&\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|\\ &\leqslant\|\mathbf{u}\|_{2}^{(2s-2+2d)/(2s)}p^{k\left(d-1-(s-1)/d\right)/(4s)}T^{o(1)}\\ &=\|\mathbf{u}\|_{2}^{(s-1+d)/s}N^{-\left((s-1)/d+1-d\right)/(2s)+o(1)}.\end{split}$
We note that we do not consider the case s>d(2d1)s>d(2d-1) as it never gives a better result, see Remark 4.1 below.
4.3. Concluding the proof
First employ the bound (4.5). Our goal is choose $s$ with (4.4) which maximises the saving in (4.5) given by
$\eta_{d}^{-}(s)=\frac{5/3+(s-2)/d-d}{2s}=\frac{1}{2d}-\frac{d(d-5/3)+2}{2ds}$
which is clearly add monotonically increasing function of $s$.
We choose the largest possible value of $s$,
$s_{1}^{-}=\left\lfloor d(2d-5/3)+2\right\rfloor$
to satisfy (4.3) and (4.4), for which we obtain
$\eta_{d}^{-}(s_{1}^{-})=\frac{1}{2d}-\frac{d(d-5/3)+2}{2d\left\lfloor d(2d-5/3)+2\right\rfloor}=\frac{\left\lfloor d(2d-5/3)\right\rfloor-d(d-5/3)}{2d\left\lfloor d(2d-5/3)+2\right\rfloor}.$
Similarly, the saving
$\eta_{d}^{+}(s)=\frac{d}{2s}$
in (4.7) is monotonically decreasing function of $s$. Hence we now choose the smallest possible value of $s$,
$s_{1}^{+}=\left\lceil d(2d-5/3)+2\right\rceil$
to satisfy (4.3) and (4.6), for which we obtain
$\eta_{d}^{+}(s_{1}^{+})=\frac{d}{2\left\lceil d(2d-5/3)+2\right\rceil}.$
Simple calculus shows that $\eta_{d}^{-}(s_{1}^{-})\geqslant\eta_{d}^{+}(s_{1}^{+})$ for $d\equiv 0,1\pmod{3}$ and $\eta_{d}^{-}(s_{1}^{-})<\eta_{d}^{+}(s_{1}^{+})$ for $d\equiv 2\pmod{3}$.
Now we can now use (4.9) and maximise the corresponding saving given by
$\vartheta_{d}(s_{2})=\frac{(s-1)/d+1-d}{2s}=\frac{1}{2d}-\frac{d(d-1)+1}{2ds},$
which is clearly a monotonically increasing function of $s$.
We choose
$s_{2}=d(2d-1)+1,$
for which we obtain
$\vartheta_{d}(s_{2})=\frac{1}{2d}-\frac{d^{2}-d+1}{2d(2d^{2}-d+1)}=\frac{d}{2(2d^{2}-d+1)}.$
Hence, using $\eta_{d}^{\pm}(s_{1}^{\pm})$ for $d\geqslant 3$ and $\vartheta_{d}(s_{2})$ for $d=1,2$, we obtain
$\max_{\psi\in\Psi_{N}(A)}\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|\leqslant\|\mathbf{u}\|_{2}^{\xi_{d}}N^{-\kappa_{d}+o(1)},$
where $\kappa_{d}$ is given by (1.6) and
$\xi_{d}=\max\left\{\frac{s_{1}^{-}-2+2d}{s_{1}^{-}},\,\frac{s_{2}-1+d}{s_{2}},\,\frac{2d^{2}}{s^{+}_{1}}\right\}$.
The proof of Theorem 1.4 concludes since the Fourier coefficients of the functions in $C^{\infty}(\mathbb{T}^{2d})$ have a rapid decay (faster than any power of $\|\mathbf{u}\|$). For more details, see [13, 15].
Remark 4.1.
To see that the case (4.8) is the only one to consider, we note that for s1>d(2d1)s-1>d(2d-1) we have
$Q_{s}(N;\mathbf{u})\leqslant\frac{1}{p^{2kd}}\|\mathbf{u}\|_{2}^{4d^{2}}T^{2s+o(1)}.$
Now Lemma 2.1 implies that
$\max_{\psi\in\Psi_{N}(A)}\left|\langle\operatorname{T}_{N}(\mathbf{u})\psi,\psi\rangle\right|\leqslant\|\mathbf{u}\|_{2}^{2d^{2}/s}N^{-d/(2s)+o(1)}.$
One easily verifies that for s1>d(2d1)s-1>d(2d-1) we have $d/(2s)\leqslant\kappa_{d}$.
Acknowledgement
The authors are very grateful to Pär Kurlberg and Zeev Rudnick for very useful comments.
During the preparation of this work, the authors were supported in part by the Australian Research Council Grant DP230100530.